OpenStudy (anonymous):

will someone show me the steps to solve this? 5/13t = -9 (five over thirteen times t equals negative nine.)

OpenStudy (anonymous):

\[\frac{5}{13}t=-9\iff t=-9\times \frac{13}{5}\]